# On prior smoothing with discrete spatial data in the context of disease mapping

### v1

@@ Abstract
Disease mapping attempts to explain observed health event counts across areal units, typically using Markov random field models. These models rely on spatial priors to account for variation in raw relative risk or rate estimates. Spatial priors introduce some degree of smoothing, wherein, for any particular unit, empirical risk or incidence estimates are either adjusted towards a suitable mean or incorporate neighbor-based smoothing. While model explanation may be the primary focus, the literature lacks a comparison of the amount of smoothing introduced by different spatial priors. Additionally, there has been no investigation into how varying the parameters of these priors influences the resulting smoothing. This study examines seven commonly used spatial priors through both simulations and real data analyses. Using areal maps of peninsular Spain and England, we analyze smoothing effects with two datasets with associated populations at risk. We propose empirical metrics to quantify the smoothing achieved by each model and theoretical metrics to calibrate the expected extent of smoothing as a function of model parameters. We employ areal maps in order to quantitatively characterize the extent of smoothing within and across the models as well as to link the theoretical metrics to the empirical metrics.
@@ 1 Introduction
Disease mapping is part of spatial epidemiology and refers to a collection of statistical techniques developed to provide accurate estimates of the geographical distribution of disease risk or rates in sets of areal units within a particular study region. Extensive literature exists on disease mapping, covering the specification, fitting, and interpretation of various models (see for example Banerjee et al., 2015; Martínez-Beneito and Botella-Rocamora, 2019). Typically, neighbor-based modeling is employed. This approach provides the joint distribution for the areal units through neighbors, specifying a local conditional distribution such that the expected value of the spatial variable at a given unit is an average of its neighboring units. Such specifications are Gaussian Markov random fields (GMRFs, Rue and Held, 2005) in the form of conditionally autoregressive (CAR) models (Besag, 1974; Besag et al., 1991). Working at the scale of areal units, these models attempt to explain the variation in risk of incidence or mortality from a disease (or diseases) over areal units in the region. The local averaging, implicit in all of these models, introduces some degree of smoothing to the inference. That is, for any particular unit, we immediately have an empirical estimate of risk or incidence/mortality rate but our modeling smooths this empirical estimate either toward a suitable mean or incorporates some version of neighbor-based smoothing. Thus, customarily, with areal unit data, an inherent component is spatial smoothing (Banerjee et al., 2015). Hence, while model assessment may be the primary objective, what has not been studied in the literature is a comparison of the amount of smoothing introduced by a particular model choice.
@@
In this regard, in the simplest univariate spatial case, the most smoothing would occur if every area had the same intercept. The least (in fact, none) smoothing would occur if each area was fitted independently with its local maximum likelihood estimate (MLE). The various spatial specifications provide something in between – neighbor-based smoothing. Therefore, we might seek to quantify the smoothing achieved by a given model specification. One can imagine metrics to quantify the amount of smoothing associated with a particular choice. For example, White et al. (2017) propose the use of predictive mean square error or rank probability scores to measure the smoothness, while Quick et al. (2021) and Song et al. (2021) offer thoughts on measuring the informativeness/performance of various CAR distributions. In any event, smoothing does not seek to minimize a goodness of fit criterion (Duncan and Mengersen, 2020).
@@
So, the contribution of this work is an investigation, both theoretical and empirical, into the extent of smoothing achieved by a given model specification. If we adopt a particular likelihood and introduce a choice of parametric spatial prior, we can ask (i) theoretically, how smoothing is expected to vary as we vary parameters in the prior, (ii) empirically, how much smoothing do we actually achieve? Further, we can ask how the smoothing varies across a collection of neighbor-based priors? Our contribution is developed within the context of disease mapping using areal data in a Bayesian framework. Specifically, this work focuses on models that smooth rates rather than risks. We specify different spatial priors, with an emphasis on CAR spatial priors, such as iCAR, LCAR or BYM, as well as a Gaussian process (GP) prior (references and details below). Moreover, we limit the theoretical analysis of smoothing behavior to the parameters of the spatial priors. We do not consider the effects of population size (which, in practice, we could not control) or spatial structure (confining our story to first-order neighborhoods) on the amount of smoothing induced. We fix the adjacency matrix to a commonly used neighborhood matrix defined by Besag et al. (1991), considering two areas as neighbors if they share a common border. However, to investigate the behavior of the models as the number of areas increases within a spatial region, we vary the level of disaggregation, specifically within continental Spain.
@@
In this study, the Poisson-Gamma non-spatial model (Clayton and Kaldor, 1987) is helpful as a baseline to develop theoretical insight into smoothing behavior since explicit calculation is available. However, to address the aforementioned questions for the models of interest, we employ both simulation studies and real data analysis. In particular, we conduct two simulation studies with different aims. The first simulation study focuses on formally quantifying the extent of smoothing and the factors influencing the level of expected smoothing for each choice of spatial prior under consideration. Additionally, it investigates whether model smoothing comparisons using theoretical metrics align with those using empirical metrics. Therefore, this simulation study performs an analysis within each spatial prior. The second simulation study compares the extent of smoothing across the proposed priors under different simulated scenarios and two different spatial structures with varying areal units and their associated neighbor structures. Then, we analyze two real datasets corresponding to lung cancer mortality counts and the corresponding population at risk for two different spatial regions, peninsular Spain and England (shown in Figure 1). We define informative and non-informative prior distributions for the parameters of the spatial priors and investigate the amount of smoothing induced both across and within spatial priors.
@@
The format of the paper is as follows. In section 2, we provide a description of the proposed Bayesian model and spatial priors used to quantify and compare the extent of smoothing. Additionally, we extract the theoretical metric proposed for analyzing the factors influencing the extent of smoothing for each spatial prior, and we present the empirical smoothing metrics used throughout the study. section 3 presents the theoretical expected smoothing for the “baseline” Poisson-Gamma model. In section 4, we present the two simulation studies: one within priors and one across priors. The real data analysis is presented in section 5. Lastly, in section 6, we summarize our contributions along with possible future investigation.
@@ 2 Modeling and Methods
In this section, the Poisson-logitNormal provides the likelihood for hierarchical modeling for disease rates. In subsection 2.1, we provide details with the collection of priors we consider. We propose a general theoretical criterion, as a function of the prior parameters to calibrate how much smoothing is expected to be induced. The empirical smoothing metrics are presented in subsection 2.2.
@@ 2.1 The Poisson-logitNormal model
With discrete spatial data, it is customary to model spatial dependence through a choice of neighbor weight matrix which provides the precision matrix, $\mathbf{Q}$, equivalently, through the conditional variances. $\mathbf{Q}$ is not nonsingular for some of the prior models; its inverse does not exist. So, notationally, we assume a normal distribution for the logit rates, that is, $\textrm{logit}\left(\mathbf{r}\right)\sim N\left(\alpha\mathbf{1_{A}},\sigma^{2}\mathbf{Q}^{-}\right)$ with $\mathbf{r}$ the vector of $r_{i}$’s, $\mathbf{1_{A}}$ a column vector of ones of size $A$, $\mathbf{Q}^{-}$ the spatial correlation matrix and - denoting the Moore-Penrose generalized inverse. When $\mathbf{Q}$ is full rank, we replace $\mathbf{Q}^{-}$ with $\mathbf{Q}^{-1}$. Therefore, we have three objects to look at here: (i) $\sigma^{2}\mathbf{Q}^{-}$, the covariance matrix of the prior defined (if it exists), (ii) $\frac{1}{\sigma^{2}}\mathbf{Q}$, the precision matrix of the prior (which always exists), and (iii) var$(r_{i}|\{r_{j},j\neq i\})=\frac{\sigma^{2}}{Q_{ii}}$, the conditional variance for $r_{i}$ (or $\kappa_{i}$) where $Q_{ii}$ denotes the $i$th diagonal element of $\mathbf{Q}$. The resulting conditional variances are smaller than the marginal variances, reflecting neighbor-based smoothing.
Analyzing the smoothing induced by different spatial priors suggests investigation of the conditional variances for the various choices of $\mathbf{Q}$, i.e., the $Q_{ii}$ through var$(r_{i}|\{r_{j},j\neq i\})=\frac{\sigma^{2}}{Q_{ii}}$ as a function of model parameters. Smaller conditional variance implies more smoothing. If we sum the conditional variances across the set of areal units, $\sigma^{2}\sum_{i}\frac{1}{Q_{ii}}$, we obtain a proposed metric for the overall amount of smoothing. In what follows, we refer to this metric as the total conditional variance (TCV). In the simulation study, for any case, we can directly calculate this metric, sometimes as an explicit parametric function. With real data, we can obtain the posterior distribution of this metric under a given prior. Further, smaller TCV implies more smoothing. We see that the magnitude of $\sigma^{2}$ will always exercise control over the amount of expected smoothing.
@@
For the non-spatial or independence prior case, suppose we place a logit normal prior on $r_{i}$, i.e., $\textrm{logit}\left(r_{i}\right)\sim N(\mu,\sigma^{2})$. Following notation from Equation 1, we view $\mu=\alpha$ and the $\kappa_{i}$ are i.i.d $N(0,\sigma^{2})$. This situation is analogous to the Poisson-Gamma setting in section 3 below. The Bayesian smoothing is toward the mean and this is a baseline for comparison. Further, the $\mathbf{TCV}=\sigma^{2}A$; we have only $\sigma^{2}$ to control smoothing.
@@
  • iCAR: $\mathbf{Q}=\mathbf{D}-\mathbf{W}$, where $\mathbf{W}$ is the spatial proximity matrix defined as $w_{ii}=0$ and $w_{ij}=1$ if the geographical units $i$ and $j$ are neighbors and 0 otherwise, $\mathbf{D}$ is a diagonal matrix whose elements are the number of neighbors of the $i$th area, $\mathbf{D}\equiv diag(w_{1}^{+};\dots;w_{A}^{+})$. Therefore, var$(r_{i}|\{r_{j},j\neq i\})=\frac{\sigma^{2}}{Q_{ii}}=\frac{\sigma^{2}}{w_{i}^{+}}$ and $\mathbf{TCV}=\sum_{i}\frac{\sigma^{2}}{w_{i}^{+}}$. Again, $\sigma^{2}$ exerts control over expect smoothing. As noted above, a comparison which we do not explore here is say, between a specification with first order neighbors vs. one with first and second order neighbors.
@@
  • BYM: This model is defined using the sum of two random effects, $\boldsymbol{\kappa}=\mathbf{u}+\mathbf{v}$ where $\mathbf{u}$ follows an iCAR prior and $\mathbf{v}$ follows an independent prior. With an iCAR prior on $u_{i}$, there is no proper distribution for $\mathbf{u}$. However, adding $\mathbf{v}$, $\boldsymbol{\kappa}$ has a proper distribution, i.e., $\boldsymbol{\kappa}\sim N\left(\boldsymbol{0},\sigma^{2}(D-W)^{-}+\tau^{2}I\right)=N\left(\boldsymbol{0},\sigma^{2}\left((D-W)^{-}+\frac{\tau^{2}}{\sigma^{2}}I\right)\right)=N\left(\boldsymbol{0},\sigma^{2}\left((D-W)^{-}+\nu I\right)\right)$ where $\nu\equiv\frac{\tau^{2}}{\sigma^{2}}$. Thus, the structure matrix is $((D-W)^{-}+\nu I))^{-1}$. Therefore, the conditional variance is $\frac{\sigma^{2}}{((D-W)^{-}+\nu I)_{ii}^{-1}}$. Again, the smaller the $\sigma^{2}$, the more smoothing. Further, it seems that $\sigma^{2}$ and $\nu$ behave reciprocally. Increasing $\nu$, i.e., $\tau^{2}$ relative to $\sigma^{2}$ will decrease smoothing, as intuition suggests. We have no closed form but $\textbf{TCV}=\sum_{i}\frac{\sigma^{2}}{((D-W)^{-}+\nu I)_{ii}^{-1}}$ can be calculated in the model fitting.
@@
  • p-CAR: $\mathbf{Q}=\mathbf{D}-\eta\mathbf{W}$, where $\eta$ is the spatial dependence parameter and $1/\epsilon_{min}<\eta<1/\epsilon_{max}$ defines a proper distribution (see Jin et al., 2007, for more details). $\epsilon_{min}$ and $\epsilon_{max}$ are the minimum and maximum eigenvalues of $\mathbf{D}^{-1/2}\mathbf{W}\mathbf{D}^{-1/2}$. For this prior, $\frac{\sigma^{2}}{Q_{ii}}=\frac{\sigma^{2}}{w_{i}^{+}}$, then $\textbf{TCV}=\sum_{i}\frac{\sigma^{2}}{w_{i}^{+}}$. As with iCAR, smoothing only depends upon $\sigma^{2}$; $\eta$ is not expected to have any impact.
  • LCAR: $\mathbf{Q}=\lambda\mathbf{R}+(1-\lambda)\mathbf{I_{A}}$, where $\mathbf{R}=\mathbf{D}-\mathbf{W}$ and $\lambda\in[0,1]$ is a spatial dependence parameter. $\lambda$ weights the structured and unstructured spatial components in $\mathbf{Q}$. This specification yields the independence case if $\lambda=0$, and iCAR if $\lambda=1$. For the LCAR prior, var$(r_{i}|\{r_{j},j\neq i\})=\frac{\sigma^{2}}{Q_{ii}}=\frac{\sigma^{2}}{\lambda w_{i+}+(1-\lambda)}=\frac{\sigma^{2}}{\lambda(w_{i+}-1)+1}$. Therefore, $\textbf{TCV}=\sum_{i}\frac{\sigma^{2}}{\lambda(w_{i+}-1)+1}$. Again, we see that as $\sigma^{2}$ decreases we have more smoothing, but also, vice versa for $\lambda$. Further, we see directly the relationship between $\sigma^{2}$ and $\lambda$. To maintain similar expected smoothing, if one goes up, the other must go down.
@@
  • BYM2: The structure matrix is an adaptation of the reparameterization of the BYM model introduced by Dean et al. (2001). It is defined as $\mathbf{Q}=\left(\lambda\mathbf{R_{*}}^{-}+(1-\lambda)\mathbf{I_{A}}\right)^{-}$, where $\mathbf{R_{*}}$ is the scaled spatial neighborhood matrix (see Riebler et al., 2016) and $\lambda\in[0,1)$. The BYM2 spatial prior has a scaled spatially structured component and its variance matrix represents a weighted average of the variances of the structured and unstructured components, which facilitates the interpretation of the model. In this case the conditional variance is $\frac{\sigma^{2}}{Q_{ii}}=\frac{\sigma^{2}}{(\lambda R^{+}+(1-\lambda)I)_{ii}^{-1}}$ and $\textbf{TCV}=\sum_{i}\frac{\sigma^{2}}{(\lambda R^{+}+(1-\lambda)I)_{ii}^{-1}}$. We can not obtain the inverse explicitly but we can calculate it for any example. The behavior in $\sigma^{2}$ is as above.
@@
Finally, we consider a spatial prior based on a geostatistical model. A geostatistical model operates at point level and, in simplest form, specifies dependence between locations as a decreasing function of distance between them. The distributional specification is customarily through a Gaussian process (GP) with a suitable choice of covariance function. Since we are analyzing discrete spatial data, i.e., a finite number of areal units, to use a GP, we supply a finite set of locations as the centroids of each of the areal units within the study area. Then, the joint probability of distribution of $\kappa_{1},\dots,\kappa_{A}$ is multivariate Normal, i.e, $\boldsymbol{\kappa}\sim N(\mathbf{0},\sigma^{2}\mathbf{R(\theta)})$ for correlation function $\rho(s_{i}-s_{j};\boldsymbol{\theta})$. We use an exponential choice, $\rho(s_{i}-s_{j};\boldsymbol{\theta})=e^{-||s_{i}-s_{j}||/\psi}$, where $\mid\mid s_{i}-s_{j}\mid\mid$ is the Euclidean distance between centroids $s_{i}$ and $s_{j}$ and $\psi$ is a range parameter. That is, for a given distance, the larger the value of $\psi$ the stronger the spatial correlation.
@@
GP $\sum_{i}\frac{\sigma^{2}}{Q_{ii}}$ $\sigma^{2}$, $\rho$
@@
BYM $\sum_{i}\frac{\sigma^{2}}{((D-W)^{-}+\nu I)_{ii}^{-1}}$ $\sigma^{2}$, $\nu=\frac{\tau^{2}}{\sigma^{2}}$
@@
BYM2 $\sum_{i}\frac{\sigma^{2}}{(\lambda R_{*}^{-}+(1-\lambda)I_{A})_{ii}^{-1}}$ $\sigma^{2}$, $\lambda$
The conditional variances for the GP will be var$(r_{i}|\{r_{j},j\neq i\})=\frac{\sigma^{2}}{Q_{ii}}$ and $\mathbf{TCV}=\sum_{i}\frac{\sigma^{2}}{Q_{ii}}$. Behavior in $\sigma^{2}$ will be as above. However, the conditional variance will depend on the set of intersite distances so is not available analytically (but, of course, these variances can be investigated in the model fitting). If we consider the case of just $A=2$ sites with an exponential correlation function, the conditional variance is $\sigma^{2}(1-e^{-2\mid\mid s_{i}-s_{j}\mid\mid/\psi})$. The suggestion is that as $\psi$ decreases, we get closer to independence, that is, we expect less smoothing. So, in our study below we consider the effect of both $\sigma^{2}$ and $\psi$ with regard to smoothing. The TCVs for the different spatial priors defined in this section are summarized in Table 1.

### v2

@@ Abstract
Disease mapping attempts to explain observed health event counts across areal units, typically using Markov random field models. These models rely on spatial priors to account for variation in raw relative risk or rate estimates. Spatial priors introduce some degree of smoothing, wherein, for any particular unit, empirical risk or incidence estimates are either adjusted towards a suitable mean or incorporate neighbor-based smoothing. While model explanation may be the primary focus, the literature lacks a comparison of the amount of smoothing introduced by different spatial priors. Additionally, there has been no investigation into how varying the parameters of these priors influences the resulting smoothing. This study examines seven commonly used spatial priors through both simulations and real data analyses. Using areal maps of peninsular Spain and England, we analyze smoothing effects using two datasets with associated populations at risk. We propose empirical metrics to quantify the smoothing achieved by each model and theoretical metrics to calibrate the expected extent of smoothing as a function of model parameters. We employ areal maps in order to quantitatively characterize the extent of smoothing within and across the models as well as to link the theoretical metrics to the empirical metrics.
@@ 1 Introduction
Disease mapping is part of spatial epidemiology and refers to a collection of statistical techniques developed to provide accurate estimates of the geographical distribution of disease risk or rates in sets of areal units within a particular study region. Extensive literature exists on disease mapping, covering the specification, fitting, and interpretation of various models (see for example Banerjee et al., 2015, Mart´ınez-Beneito and Botella-Rocamora, 2019). Typically, neighbor-based modeling is employed. This approach provides the joint distribution for the areal units through neighbors, specifying a local conditional distribution such that the expected value of the spatial variable at a given unit is an average of its neighboring units. Such specifications are Gaussian Markov random fields (GMRFs, Rue and Held, 2005) in the form of conditionally autoregressive (CAR) models (Besag, 1974, Besag et al., 1991). Working at the scale of areal units, these models attempt to explain the variation in risk of incidence or mortality from a disease (or diseases) over areal units in the region. The local averaging, implicit in all of these models, introduces some degree of smoothing to the inference. That is, for any particular unit, we immediately have an empirical estimate of risk or incidence/mortality rate but our modeling smooths this empirical estimate either toward a suitable mean or incorporates some version of neighbor-based smoothing. Thus, customarily, with areal unit data, an inherent component is spatial smoothing (Banerjee et al., 2015). However, the degree of smoothing must be carefully managed. Insufficient smoothing results in noisy maps dominated by random variation, while excessive smoothing can obscure the detection of both high- and low-risk areas. Since one of the main goals of disease mapping is to identify spatial patterns that inform public health decisions, it is crucial to understand and assess the extent of smoothing induced by the chosen model to ensure valid and actionable conclusions. Hence, while model assessment is often emphasized and may be the primary objective, the amount of smoothing introduced by a particular model choice has not been studied in the literature. This paper aims to address this gap.
@@
In this regard, in the simplest univariate spatial case, with no risk factors, the most smoothing would occur if every area had the same intercept. The least (in fact, none) smoothing would occur if each area was fitted independently with its local maximum likelihood estimate (MLE). The various spatial specifications provide something in between – neighbor-based smoothing. Therefore, we might seek to quantify the smoothing achieved by a given model specification. One can imagine metrics to quantify the amount of smoothing associated with a particular choice. For example, White et al. (2017) propose the use of predictive mean square error or rank probability scores to measure the smoothness, while Quick et al. (2021) and Song et al. (2021) offer thoughts on measuring the informativeness/performance of various CAR distributions. In any event, smoothing does not seek to minimize a goodness of fit criterion (Duncan and Mengersen, 2020).
@@
So, the contribution of this work is an investigation, both theoretical and empirical, into the extent of smoothing achieved by a given model specification. If we adopt a particular likelihood and introduce a choice of parametric spatial prior, we can ask (i) theoretically, how smoothing is expected to vary as we vary parameters in the prior, (ii) empirically, how much smoothing do we actually achieve? Further, we can ask how the smoothing varies across a collection of neighbor-based priors? Our contribution is developed within the context of disease mapping using areal data in a Bayesian framework. Specifically, this work focuses on models that smooth rates rather than risks. We specify different spatial priors, with an emphasis on CAR spatial priors, such as iCAR, LCAR or BYM, as well as a Gaussian process (GP) prior (references and details below). Moreover, we limit the theoretical analysis of smoothing behavior to the parameters of the spatial priors. We do not consider the effects of population size (which, in practice, we could not control) or spatial structure (confining our story to first-order neighborhoods) on the amount of smoothing induced. We fix the adjacency matrix to a commonly used neighborhood matrix defined by Besag et al. (1991), considering two areas as neighbors if they share a common border. However, to investigate the behavior of the models as the number of areas increases within a spatial region, we vary the level of disaggregation, specifically within peninsular Spain. In summary, our contribution offers novelty in illuminating theoretical smoothing behavior with regard to prior parameters across a collection of priors. We see potential practical utility in terms of comparative expected smoothness across these priors after model fitting, anticipating practitioner decision-making between less smoothed maps which are noisier or more smoothed maps which obscure risk detection. We illustrate this issue in section 5.
@@
In this study, the Poisson-Gamma non-spatial model (Clayton and Kaldor, 1987) is helpful as a baseline to develop theoretical insight into smoothing behavior since explicit calculation is available. However, to address the aforementioned questions for the models of interest, we employ both simulation studies and real data analyses. In particular, we conduct two simulation studies with different aims. The first simulation study focuses on formally quantifying the extent of smoothing and the factors influencing the level of expected smoothing for each choice of spatial prior under consideration. Additionally, it investigates whether model smoothing comparisons using theoretical metrics align with those using empirical metrics. Therefore, this simulation study performs an analysis within each spatial prior. The second simulation study compares the extent of smoothing across the proposed priors under different simulated scenarios and two different spatial structures with varying areal units and their associated neighbor structures. Then, we analyze two real datasets corresponding to lung cancer mortality counts and the corresponding population at risk for two different spatial regions, peninsular Spain and England. We define informative and non-informative prior distributions for the parameters of the spatial priors and investigate the amount of smoothing induced both across and within spatial priors.
@@
The format of the paper is as follows. In section 2, we provide a description of the employed Bayesian model and spatial priors used to quantify and compare the extent of smoothing. Additionally, we derive the theoretical metric proposed for analyzing the factors influencing the extent of smoothing for each spatial prior, and we present the empirical smoothing metrics used throughout the study. section 3 presents the theoretical expected smoothing for the “baseline” Poisson-Gamma model. In section 4, we present the two simulation studies: one within priors and one across priors. The real data analysis is presented in section 5. Lastly, in section 6, we summarize our contributions along with possible future investigation.
@@ 2 Modeling and Methods
In this section, the Poisson-logitNormal provides the likelihood for hierarchical modeling of disease rates. In subsection 2.1, we provide details with the collection of priors we consider. We propose a general theoretical criterion, as a function of the prior parameters to calibrate how much smoothing is expected to be induced. The empirical smoothing metrics are presented in subsection 2.2.
@@ 2.1 The Poisson-logitNormal model
[…]
$\text{TCV}=\sum_{i=1}^{n}\operatorname{var}(r_{i}\mid\{r_{j},\,j\neq i\})=\sum_{i=1}^{n}\frac{\sigma^{2}}{Q_{ii}}.$
This metric serves as a scalar summary of the smoothing implied by a given prior. In the simulation study, for any case, we can directly calculate this metric, sometimes as an explicit parametric function. With real data, we can obtain the posterior distribution of this metric under a given prior. Further, smaller TCV implies more smoothing. We see that the magnitude of $\sigma^{2}$ will always exercise control over the amount of expected smoothing.
@@
For the non-spatial or independence prior case, suppose we place a logit normal prior on $r_{i}$, i.e., $\textrm{logit}\left(r_{i}\right)\sim N(\mu,\sigma^{2})$. Following notation from Equation 1, we view $\mu=\alpha$ and the $\kappa_{i}$ are i.i.d $N(0,\sigma^{2})$. This situation is analogous to the Poisson-Gamma setting in section 3 below. The Bayesian smoothing is toward the mean and this is a baseline for comparison. Further, the $\textrm{TCV}=\sigma^{2}A$; we have only $\sigma^{2}$ to control smoothing.
@@
[…]
@@
  • BYM: This model is defined using the sum of two random effects, $\boldsymbol{\kappa}=\mathbf{u}+\mathbf{v}$ where $\mathbf{u}$ follows an iCAR prior and $\mathbf{v}$ follows an independent prior. With an iCAR prior on $u_{i}$, there is no proper distribution for $\mathbf{u}$. However, adding $\mathbf{v}$, $\boldsymbol{\kappa}$ has a proper distribution, i.e., $\boldsymbol{\kappa}\sim N\left(\boldsymbol{0},\sigma^{2}(\mathbf{D}-\mathbf{W})^{-}+\tau^{2}\mathbf{I_{A}}\right)=N\left(\boldsymbol{0},\sigma^{2}\left((\mathbf{D}-\mathbf{W})^{-}+\frac{\tau^{2}}{\sigma^{2}}\mathbf{I_{A}}\right)\right)=N\left(\boldsymbol{0},\sigma^{2}\left((\mathbf{D}-\mathbf{W})^{-}+\nu\mathbf{I_{A}}\right)\right)$ where $\nu\equiv\frac{\tau^{2}}{\sigma^{2}}$. Thus, the structure matrix is $((\mathbf{D}-\mathbf{W})^{-}+\nu\mathbf{I_{A}}))^{-1}$. Therefore, the conditional variance is $\frac{\sigma^{2}}{((\mathbf{D}-\mathbf{W})^{-}+\nu\mathbf{I_{A}})_{ii}^{-1}}$. Again, the smaller the $\sigma^{2}$, the more smoothing. Further, it seems that $\sigma^{2}$ and $\nu$ behave reciprocally. Increasing $\nu$, i.e., $\tau^{2}$ relative to $\sigma^{2}$ will decrease smoothing, as intuition suggests. We have no closed form but $\textrm{TCV}=\sum_{i}\frac{\sigma^{2}}{((\mathbf{D}-\mathbf{W})^{-}+\nu\mathbf{I_{A}})_{ii}^{-1}}$ can be calculated in the model fitting.
@@
  • p-CAR: $\mathbf{Q}=\mathbf{D}-\eta\mathbf{W}$, where $\eta$ is the spatial dependence parameter and $1/\epsilon_{min}<\eta<1/\epsilon_{max}$ defines a proper distribution (see Jin et al., 2007, for more details). $\epsilon_{min}$ and $\epsilon_{max}$ are the minimum and maximum eigenvalues of $\mathbf{D}^{-1/2}\mathbf{W}\mathbf{D}^{-1/2}$. For this prior, $\frac{\sigma^{2}}{Q_{ii}}=\frac{\sigma^{2}}{w_{i}^{+}}$, then $\textrm{TCV}=\sum_{i}\frac{\sigma^{2}}{w_{i}^{+}}$. As with iCAR, smoothing only depends upon $\sigma^{2}$; $\eta$ is not expected to have any impact.
[…]
@@
  • BYM2: The structure matrix is an adaptation of the reparameterization of the BYM model introduced by Dean et al. (2001). It is defined as $\mathbf{Q}=\left(\lambda\mathbf{R_{*}}^{-}+(1-\lambda)\mathbf{I_{A}}\right)^{-}$, where $\mathbf{R_{*}}$ denotes the spatial neighborhood matrix from the iCAR model, $\mathbf{R}=\mathbf{D}-\mathbf{W}$, scaled according to the geometric mean of the marginal variance (see Riebler et al., 2016, for further details) and $\lambda\in[0,1]$. The BYM2 spatial prior has a scaled spatially structured component and its variance matrix represents a weighted average of the variances of the structured and unstructured components, which facilitates the interpretation of the model. In this case the conditional variance is $\frac{\sigma^{2}}{Q_{ii}}=\frac{\sigma^{2}}{(\lambda\mathbf{R_{*}}^{-}+(1-\lambda)\mathbf{I_{A}})_{ii}^{-1}}$ and $\textrm{TCV}=\sum_{i}\frac{\sigma^{2}}{(\lambda\mathbf{R_{*}}^{-}+(1-\lambda)\mathbf{I_{A}})_{ii}^{-1}}$. We can not obtain the inverse explicitly but we can calculate it for any example. The behavior in $\sigma^{2}$ is as above.
@@
Finally, we consider a spatial prior based on a geostatistical model. A geostatistical model operates at point level and, in simplest form, specifies dependence between locations as a decreasing function of distance between them. The distributional specification is customarily through a Gaussian process (GP) with a suitable choice of covariance function. Since we are analyzing discrete spatial data, i.e., a finite number of areal units, to use a GP, we supply a finite set of locations as the centroids of each of the areal units within the study area. Then, the joint probability of distribution of $\kappa_{1},\dots,\kappa_{A}$ is multivariate Normal, i.e, $\boldsymbol{\kappa}\sim N(\mathbf{0},\sigma^{2}\mathbf{R(\theta)})$ where each element of the covariance matrix $\mathbf{R(\theta)}$ is specified by a correlation function $\rho(s_{i}-s_{j};\boldsymbol{\theta})$. We illustrate with an exponential choice, $\rho(s_{i}-s_{j};\boldsymbol{\theta})=e^{-||s_{i}-s_{j}||/\psi}$, where $\mid\mid s_{i}-s_{j}\mid\mid$ is the Euclidean distance between centroids $s_{i}$ and $s_{j}$ and $\psi$ is a range parameter. That is, for a given distance, the larger the value of $\psi$ the stronger the spatial correlation.
@@
GP $\sum_{i}\frac{\sigma^{2}}{Q_{ii}}$ $\sigma^{2}$, $\psi$
@@
BYM $\sum_{i}\frac{\sigma^{2}}{((\mathbf{D}-\mathbf{W})^{-}+\nu\mathbf{I_{A}})_{ii}^{-1}}$ $\sigma^{2}$, $\nu=\frac{\tau^{2}}{\sigma^{2}}$
@@
BYM2 $\sum_{i}\frac{\sigma^{2}}{(\lambda\mathbf{R_{*}}^{-}+(1-\lambda)\mathbf{I_{A}})_{ii}^{-1}}$ $\sigma^{2}$, $\lambda$
[…]